After a brief look this morning, I called it and declared that we were ok to close up. The access connector is almost all buttoned up, and both ETM doors are on.
Basically nothing moved since last night, which is good. Jenne and I were a little bit worried about how the input pointing might have been effected by our moving of the green periscope in the MC chamber.
First thing this morning I went into the BS chamber to check out the alignment situation there. I put the targets on the PRM and BS cages. We were basically clear through the PRM aperture, and in retro-reflection.
The BS was not quite so clear. There is a little bit of clipping through the exit aperture on the X arm side. However, it didn't seem to me like it was enough to warrant retouching all the input alignment again, as that would have set us back another couple of days at least.
Both arm green beams are cleaning coming out, and are nicely overlapping with the IR beams at the BS (we even have a clean ~04 mode from the Y arm). The AS and REFL spots look good. IPANG and IPPOS are centered and haven't moved much since last night. We're ready to go.
The rest of the vertex doors will go on after lunch.